Hickory Company manufactures two products—13,000 units of Product Y and 5,000 units of Product Z. The company uses a plantwide overhead rate based on direct labor-hours. It is considering implementing an activity-based costing (ABC) system that allocates all $745,200 of its manufacturing overhead to four cost pools. The following additional information is available for the company as a whole and for Products Y and Z:

Activity Cost Pool Activity Measure Estimated Overhead Cost Expected Activity
Machining Machine-hours $ 242,400 12,000 MHs
Machine setups Number of setups $ 114,400 220 setups
Product design Number of products $ 86,000 2 products
General factory Direct labor-hours $ 302,400 14,400 DLHs
Activity Measure Product Y Product Z
Machine-hours 7,200 4,800
Number of setups 40 180
Number of products 1 1
Direct labor-hours 8,200 6,200

9. Using the ABC system, how much total manufacturing overhead cost would be assigned to Product Y?
10.Using the ABC system, how much total manufacturing overhead cost would be assigned to Product Z?
11. Using the plantwide overhead rate, what percentage of the total overhead cost is allocated to Product Y and Product Z?
12. Using the ABC system, what percentage of the Machining costs is assigned to Product Y and Product Z?
13. Using the ABC system, what percentage of Machine Setups cost is assigned to Product Y and Product Z?
14. Using the ABC system, what percentage of the Product Design cost is assigned to Product Y and Product Z?
15. Using the ABC system, what percentage of the General Factory cost is assigned to Product Y and Product Z?
Answer #1

9) Total manufacturing overhead cost assigned to product Y = (20.2*7200+520*40+43000+21*8200) = $381440

10) Total manufacturing overhead cost assigned to product Z = (20.2*4800*520*180+43000+21*6200) = $363760

11) Percentage

Product Y Product Z
Direct labor hour 8200/14400 = 57% 6200/14400 = 43%

12) Percentage

Product Y Product Z
Machine hour 7200/12000 = 60% 40%

13) Percentage

Product Y Product z
Machine setup 40/220 = 18.18% 180/220 = 81.82%

14) Percentage

Product Y Product z
Product design 1/2 = 50% 50%

15) Percentage

Product Y Product z
General factory 8200/14400 = 57% 43%